OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 4x2 + 28x + 49 square feet, what is the length of one side of the room? (2x + 7) feet (7x - 2) feet (7x + 2) feet (2x - 7) feet

OpenStudy (anonymous):

3.5!!

OpenStudy (anonymous):

4x^2+28x+49 = (2x+7)^2 Since the room is square, its area formula is A = s^2, where s = length of one side so the length of one side = the square root of the equation, which = (2x+7) feet

OpenStudy (anonymous):

AH YE right thenx